OpenStudy (anonymous):

Let f be a function defined by f(x)=(x/5)-(1/5). Determine the value of x for which f(x)= -4

OpenStudy (trojanpoem):

when f(x) = -4 -4 = x/5 - 1/5 -4 + 1/5 = x/5 -20/5 + 1/5 = x/5 -19/5 = x/5 x= -19

OpenStudy (anonymous):

thank you very much.